P is a point on the parabola whose ordinate

equals its abscissa. A normal is drawn to the parabola at

P to meet it again at Q. If S is the focus of the parabola

then the product of the slopes of SP and SQ is-



1
Expert's answer
2021-09-21T17:41:28-0400

Let point P(at2, 2at) lie on parabola y2=4ax

With ordinate of P equals to its abscissa,

at2=2at. Solving, t=2

P is (4a,4a)

Equation of normal to the parabola y2=4ax at (x1,y1) is

y-y1= -y1(x-x1)/2a

At P(4a, 4a), equation of normal simplifies to

y+2x = 12a

Given also that y2= 4ax, and solving the equations, y=4a and x= 4a or y= -6a and x=9a

Hence S(a,0) P(4a, 4a) Q(9a, -6a) gives slope

SP = 4/3 and slope SQ = -6/8

Product = 4/3 *-6/8 = -1